Hello professor. I'm having trouble with linear function problems.
Suppose the sales of a particular brand of appliance satisfy the linear model y=90x + 3400, where "y" represents the number of sales in year "x", with x= 0 corresponding to 1982. Find the number of sales in 1990.
(A) 4120 (B) 8150 (C) 8240 (D) 4030
Thank you!!!

Answer by Earlsdon(6294)

The key here is to find the value of x.
If x=0 corresponds to 1982, 1990 is 1982+8 so x = 8 for 1990. Substitute x=8 into the given equation.
y = 90(8)+3400 Evaluate.
y = 720+3400 Add.
y = 4120